# LG-ANNA-Embedding Technical Report

### v1

@@ 1 Introduction
In this work, we present LG-ANNA-Embedding, a generalized embedding model built upon Mistral-7B and enhanced with in-context learning, soft supervision, and adaptive hard-negative sampling. Without modifying the underlying model architecture, our approach achieves strong generalization across a wide range of retrieval and non-retrieval tasks. Evaluated on the newly introduced MTEB(English, v2) benchmark [5], our model demonstrates competitive performance across 41 diverse tasks, ranking among the top systems in terms of Borda score. Our findings highlight the effectiveness of combining in-context LLM capabilities with high-quality training signals for scalable and accurate embedding generation.
@@ 4.3 Adaptive Margin-Based Mining Strategies
LG-ANNA-Embedding(Ours) 66.18 49.13 59.25 88.67 89.97 86.69 38.93 74.12
@@ 5.2 MTEB Results
Compared to existing top-ranked models on the MTEB English (v2) benchmark, our proposed LG-ANNA-Embedding demonstrates consistently strong performance across a broad range of task categories. Notably, it achieves high scores in both retrieval (66.18; 2nd place), sts (86.69; 2nd place), pair-classification (88.67; 2nd place) and summarization (38.93; 1st place), ranking among the top models in these categories. In particular, the model yields the highest summarization score across all evaluated systems, outperforming both larger commercial models such as gemini-embedding-001 (38.28) and Seed1.5-Embedding (36.44). Although LG-ANNA-Embedding reports a slightly lower mean score (74.12) compared to Seed1.5-Embedding (74.76), it surpasses it in overall Borda ranking, reflecting stronger and more consistent performance across diverse task types. These results suggest that our model strikes a favorable balance between generalization and task-specific performance, particularly excelling in information retrieval and summarization.
@@ 6 Conclusion
In this report, we presented LG-ANNA-Embedding, a unified and instruction-based framework for generating high-quality, generalized text embeddings using decoder-only LLMs. By combining in-context learning with soft supervision from a high-performance retriever-reranker pipeline and adaptive margin-based hard-negative mining, our method effectively leverages the inherent generalization capabilities of large language models without requiring full model fine-tuning or architectural changes.
@@ B  Model License
LG-ANNA-Embedding Model License: Apache 2.0 and CC BY-NC 4.0
This License Agreement (“Agreement”) is entered into between you (“Licensee”) and LG Management Development Institute Co., Ltd. (“Licensor”), governing the use of the LG-ANNA-Embedding model (“Model”). By downloading, installing, copying, or using the Model, you agree to comply with and be bound by the terms of this Agreement. If you do not agree to all the terms, you must not download, install, copy, or use the Model. This Agreement constitutes a binding legal agreement between the Licensee and Licensor.
The LG-ANNA-Embedding model is released under the Apache License 2.0 for software components and the Creative Commons Attribution-NonCommercial 4.0 International License (CC BY-NC 4.0) for data and model weights. For details, please refer to the full license texts at:

### v2

# LGAI-Embedding-Preview Technical Report

@@ 1 Introduction
In this work, we present LGAI-Embedding-Preview, a generalized embedding model built upon Mistral-7B and enhanced with in-context learning, soft supervision, and adaptive hard-negative sampling. Without modifying the underlying model architecture, our approach achieves strong generalization across a wide range of retrieval and non-retrieval tasks. Evaluated on the newly introduced MTEB(English, v2) benchmark [5], our model demonstrates competitive performance across 41 diverse tasks, ranking among the top systems in terms of Borda score. Our findings highlight the effectiveness of combining in-context LLM capabilities with high-quality training signals for scalable and accurate embedding generation.
@@ 4.3 Adaptive Margin-Based Mining Strategies
LGAI-Embedding-Preview(Ours) 66.18 49.13 59.25 88.67 89.97 86.69 38.93 74.12
@@ 5.2 MTEB Results
Compared to existing top-ranked models on the MTEB English (v2) benchmark, our proposed LGAI-Embedding-Preview demonstrates consistently strong performance across a broad range of task categories. Notably, it achieves high scores in both retrieval (66.18; 2nd place), sts (86.69; 2nd place), pair-classification (88.67; 2nd place) and summarization (38.93; 1st place), ranking among the top models in these categories. In particular, the model yields the highest summarization score across all evaluated systems, outperforming both larger commercial models such as gemini-embedding-001 (38.28) and Seed1.5-Embedding (36.44). Although LGAI-Embedding-Preview reports a slightly lower mean score (74.12) compared to Seed1.5-Embedding (74.76), it surpasses it in overall Borda ranking, reflecting stronger and more consistent performance across diverse task types. These results suggest that our model strikes a favorable balance between generalization and task-specific performance, particularly excelling in information retrieval and summarization.
@@ 6 Conclusion
In this report, we presented LGAI-Embedding-Preview, a unified and instruction-based framework for generating high-quality, generalized text embeddings using decoder-only LLMs. By combining in-context learning with soft supervision from a high-performance retriever-reranker pipeline and adaptive margin-based hard-negative mining, our method effectively leverages the inherent generalization capabilities of large language models without requiring full model fine-tuning or architectural changes.
@@ B  Model License
LGAI-Embedding-Preview Model License: Apache 2.0 and CC BY-NC 4.0
This License Agreement (“Agreement”) is entered into between you (“Licensee”) and LG Management Development Institute Co., Ltd. (“Licensor”), governing the use of the LGAI-Embedding-Preview model (“Model”). By downloading, installing, copying, or using the Model, you agree to comply with and be bound by the terms of this Agreement. If you do not agree to all the terms, you must not download, install, copy, or use the Model. This Agreement constitutes a binding legal agreement between the Licensee and Licensor.
The LGAI-Embedding-Preview model is released under the Apache License 2.0 for software components and the Creative Commons Attribution-NonCommercial 4.0 International License (CC BY-NC 4.0) for data and model weights. For details, please refer to the full license texts at: